an investment adviser invested $14000 in two accounts. one investment earned 8% annual simple interest, an the other investment earned 6.5% annual simple interest. the amount of interest earned for 1 year was $1024. how much was investedin each account?
Answer by checkley71(8403)

.08X+.065(14000-X)=1024
.08X+910-.065X=1024
.015X=1024-910
.015X=114
X=114/.015
X=7600 INVESTED @ 8%
14000-7600=6400 INVESTED @ 6.5%
PROOF
.08*7600+.065*6400=1024
608+416=1024
1024=1024